OpenStudy (anonymous):

[Calculus] derivative using the quotient rule y=5/cosx

OpenStudy (campbell_st):

well you can write your equation as \[y = 5 \sec(x) \] or perhaps \[y = 5(\cos(x))^{-1} \] and apply the cain rule...

OpenStudy (campbell_st):

oops chain rule

OpenStudy (anonymous):

\[\prime =\frac{ 5sinx }{ \cos ^{2}x}\]

OpenStudy (anonymous):

is it right? :)

OpenStudy (cgreenwade2000):

No. Just use the quotient rule.

OpenStudy (cgreenwade2000):

Im pretty sure thats what they want you to do since it tells you to solve it that way.

OpenStudy (anonymous):

Derivative of Trig func. using quotient rule.

OpenStudy (anonymous):

y'=5/cosx = {[(cosx-sinx)(cosx)]-[(sinx+cosx)(-sinx)]}/cos^2x = (cos^2x-sinxcosx)-(-sin^2x-sinxcosx)/cos^2x =(cos^2x-sinxcosx+sin^2x+sinxcosx/cos^2x =cos^2x+sin^2x/cos^2x

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Thank you! :)